OpenStudy (anonymous):

If x and y vary directly, and x = 9 when y = 3, what will be the value of x when y = 15?

OpenStudy (anonymous):

Please show how you did it

OpenStudy (anonymous):

first divide 15(the new number) by 3( the original), and you get 5. multiply 9(the original) by 5 for the answer. I hope I did a good job explaining.

OpenStudy (nurali):

if x and y vary directly, then y = kx from x = 9 and y = 3 we have: 3 = k*9 so k = 3/9 k=1/3 and we have: y = (1/3)x evaluating at y = 15: 15 = (1/3)x x = 15*3 x = 45

OpenStudy (anonymous):

hi :DD

OpenStudy (anonymous):

It doesn't get easier than what @Nurali said

OpenStudy (anonymous):

yes thank so much

OpenStudy (nurali):

Anytime.